OpenStudy (anonymous):

Solve for x: [(x-3)/(x-1)] less tahn or equal to [(4/(x+8)]

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

last one for the night. you have to actually put everything on one side of the equal sign in one term

OpenStudy (anonymous):

\[\frac{x-3}{x-1}\leq \frac{4}{x+8}\] \[\frac{x-3}{x-1}-\frac{4}{x+8}\leq 0\] \[\frac{(x-3)(x+8)-4(x-1)}{(x-1)(x+8)}\] see why i said ick?

OpenStudy (anonymous):

\[\frac{x^2-x+20}{(x-1)(x+8)}\leq 0\] \[\frac{(x-5)(x+4)}{(x-1)(x+8)}\leq 0\]

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

zeros of these factors are -8, -4, 1, 5 so we have to break up the line into 5 pieces ----------- -8 ------------- -5 -------------- 1------------ 5 --------------- and determine the sign on each interval

OpenStudy (anonymous):

oh is this liek a number line?

OpenStudy (anonymous):

i give you a hint. it will be positive , negative, positive, negative positive

OpenStudy (anonymous):

yeah a number line broken into 5 parts at the zeros of each factor

OpenStudy (anonymous):

it changes sign at each zero, so you really only have to check one and you will know them all

OpenStudy (anonymous):

im kinda confused

OpenStudy (anonymous):

from \[(-\infty,-8)\] positive from \[(-8,-5)\] negative and so on

OpenStudy (anonymous):

yeah am sure you are

OpenStudy (anonymous):

it is all algebra up to this step \[\frac{(x-5)(x+4)}{(x-1)(x+8)}\leq 0\]

OpenStudy (anonymous):

painful annoying but necessary algebra. now you have 4 factors rigth?

OpenStudy (anonymous):

yeah i se taht...so tehn u set tehm to 0 n u got -8, -4, 1, 5. I got it up to there

OpenStudy (anonymous):

ok so you understand that \[<0\] is a synonym for negative

OpenStudy (anonymous):

so now the question is, for which values of x will this monstrosity be negative

OpenStudy (anonymous):

it will change sign at those zeros. it will go from being positive to negative or negative to positive. so your job now is to figure out over which of those intervals it is positive, over which it is negative, and pick the negative ones as our answer

OpenStudy (anonymous):

the intervals are \[(-\infty,-8),(-8,-5),(-5,1),(1,5),(5,\infty)\]

OpenStudy (anonymous):

and you have to pick the right ones. the ones for which it is negative. do you know how to tell?

OpenStudy (anonymous):

sorry not really..teh comp frozwe

OpenStudy (anonymous):

ok you can just check over any interval. pick any number. i pick 0. plug in 0 and see what you get

OpenStudy (anonymous):

you dont even have to compute

OpenStudy (anonymous):

if you replace x by 0, two factors are negative and two are positive, so it will be positive

OpenStudy (anonymous):

that means on the entire interval (-5,1) it is positive because 0 is in that interval. so it goes like this Positive Negative Positive Negative Positive respectively

OpenStudy (anonymous):

you want the negative ones, so your solution is \[(-8,-5]\cup (1,5]\]

OpenStudy (anonymous):

notice the careful placement of ( vs [ you have less than or equal to so you use [ or ] except if that number would make the DENOMINATOR 0. that is why i wrote \[(-8,-5]\] open and closed

OpenStudy (anonymous):

i see, wow thsi was a long and tedious probulem huh?

OpenStudy (anonymous):

wait so is (−8,−5]∪(1,5] teh solution or (−8,−5]

OpenStudy (anonymous):

both

OpenStudy (anonymous):

oh ok tahnks soo much...i was wrking on ths ifor awhile n i cudnt figure it out